# Turing instability and 2-D pattern formation in reaction-diffusion systems derived from kinetic theory

### v1

@@ 1 Introduction
Reaction-diffusion equations play a crucial role in various areas of applied sciences, offering a powerful framework for modeling diverse mechanisms in complex systems, that involve both interactions among the components and diffusive phenomena. The literature on this topic is extensive and includes applications in biology (e.g., epidemic spread [2, 12]), chemistry (e.g., autocatalytic reactions [22, 13]), ecology (e.g., vegetation patterns [1, 15]), medicine (e.g., tumor growth, autoimmune diseases [11, 14, 4]), and economics (e.g., markets and migration [3]).
@@
The literature on the subject typically proposes mathematical models for densities, in which macroscopic parameters are set empirically. This approach usually neglects possible microscopic dynamics between agents, focusing instead on the average effect over the entire population. It is therefore desirable to connect the macroscopic level with the microscopic one, and in this regard, the classical and extended kinetic theory proves to be a valuable tool. For the sake of completeness, it is important to underline that, from a broader perspective, kinetic theory offers the possibility of working across multiple scales, leading to a deeper understanding of collective phenomena and allowing, for example, insight into how a population is distributed around the average value; some steps in this direction have recently been taken [8, 19].
@@
In this paper, we focus on two models [6, 20] recently derived in the classical context of gasdynamics in proper multiple scale regimes; references show how linear and nonlinear diffusion terms can be obtained through a rigorous limit of Boltzmann-like kinetic equations governing a mixture of monatomic and polyatomic gases. More precisely, in the continuum limit, in [6] components densities are governed by a Brusselator-type description, where a further parameter appears in addition to those present in the classic model for autocatalytic reactions; whereas, in [20] a model with nonlinear cross diffusion is derived, reproducing the diffusive terms of [9] but differing from it in the reactive contributions. In both cases, the parameters in the macroscopic setting can be expressed in terms of the microscopic quantities, such as mass, internal energy and collision frequency.
@@
The paper is organized as follows: we recall the derivation of the two models as diffusive limit of proper kinetic equations in Section 2. The Turing instability and the pattern formation for the Brusselator-type model and the description with cross diffusion are in Section 3 and 4, respectively: the emergence of non-uniform stable solutions is investigated through weakly nonlinear analysis and, confirmed and extended numerically. Some concluding remarks are given in Section 5.
@@ 2 Kinetic equations and diffusive limits
We consider a binary gas mixture diffusing in a host medium. The first constituent $Y$ is assumed to be polyatomic with two different energy levels $E_{1}$ and $E_{2}$; therefore, it can be split in two components $Y_{1}$ and $Y_{2}$. The second gas $Z$ is supposed to be monatomic, endowed with a unique energy level $E_{Z}$. We indicate by $m_{Y}$ and $m_{Z}$ the masses of polyatomic and monatomic gas particles, respectively.
@@ 2.1 Linear diffusion
It is important to underline that, through the derivation, an explicit expression of $n_{Z}$ as quadratic function of $n_{1}$ is deduced; this algebraic relation allows to reduce the number of variables in the diffusive limit. The coefficients for macroscopic observables are positive constants defined in terms of the microscopic parameters (collision frequencies, masses and energy levels), providing some restrictions for the range of validity.
@@ 2.2 Nonlinear cross diffusion
where $N=n_{1}+n_{2}$ is the total number density of the polyatomic component and the use of an algebraic relation relating $n_{1}$, $n_{2}$ and $n_{Z}$ has been made. Once again, the parameters are positive constants, depending on microscopic parameters as provided below
@@ 3.1 Turing instability
By recalling the expressions coefficients in terms of microscopic parameters of the mixture (2.12)-(2.15), and fixing values as follows
@@
We individuate the values for energy levels $E_{2}$ and $E_{Z}$ leading to condition (3.15) and (3.16) to be satisfied, as reported in Figure 1. In particular, we can distinguish four regions. In region I, condition (3.3) is not satisfied and the spatially homogeneous equilibrium is unstable. In regions II, III, and IV, instead, (3.3) holds, and either both conditions (3.15) and (3.16), only (3.16), or neither of the conditions in (3.15) and (3.16) are satisfied, respectively. Consequently, pattern formation can arise only when taking parameters in region II. It should be emphasized that, within region II, the parameter $d$ is not restricted to the value 1, which once again highlights the broader spectrum of scenarios due to the presence of this additional parameter.
@@ 3.2 Weakly nonlinear analysis
where by c.c. we indicate the complex conjugate of the first terms and $\sigma=\dfrac{a\xi\,\sqrt{d}\,}{1+a\xi\,\sqrt{d}\,}$.
@@
$+\sum_{\begin{subarray}{c}j=1,2,3\\ l\equiv j+1\mbox{ \tiny mod }3\end{subarray}}\left(\begin{array}[]{c}X_{1}\\ Y_{1}\end{array}\right)\,\mathcal{W}_{j}\,\overline{\mathcal{W}}_{l}\,e^{\,i\,(\mathbf{k}_{j}-\mathbf{k}_{l})\cdot\mathbf{x}}+\mbox{ c.c.}.$
@@ 3.3 Numerical simulations
In this section we investigate the behavior of system (2.11) by performing numerical simulations. We simulate a two-dimensional square domain of length $L=12\pi$, discretized into a $\times100$ grid. The solution is initialized by adding a random perturbation to the equilibrium values. The diffusive terms are discretized using a second-order finite-difference method, the source terms are evaluated locally, and the solution is marched in time from the initial conditions until convergence using a first-order forward-Euler scheme. It is worth noting that once the solution is converged to a steady state, the error associated with the time discretization vanishes. Therefore, using a first-order time-integration scheme is acceptable for our purposes. The time step is selected adaptively during the simulation such that the local Von Neumann number—estimated by computing the maximum diffusion coefficient through the domain—remains below a certain threshold. At the boundary of the domain, we impose zero-gradient Neumann conditions. We do so by utilizing a layer of ghost cells and adapting their value as the simulation evolves. The method is implemented in a modified version of the open-source Hyper2D solver [7] tailored to reaction-diffusion systems.
@@
For the sake of brevity, we do not report the patterns in regions I and II, since their shape is similar to that of the patterns in regions IV and V, respectively; the only difference concerns the phase angle $\phi$, which passes from $\pi$ to 0, giving rise to the so-called reentrant hexagons [24].
@@ 4.1 Turing instability and weakly nonlinear analysis
As in the previous case, we want to focus on the microscopic features of the model, by finding suitable sets of the microscopic parameters that lead to the conditions above to be satisfied by the macroscopic coefficients, through the relations (2.19)-(2.21). For our purposes, we fix some values as follows
@@ 4.2 Numerical simulations
We point out that, as shown also in the one-dimensional case in [20], the numerical results suggest that the profile of $N$ is sum of more than one mode in each direction $\mathbf{k}_{i}$. Therefore, a deeper analysis of the solution near the critical threshold is needed; this will be subject of future investigation.
@@ 5 Conclusions
The proposed analysis shows the strength of these tools in linking macroscopic dynamics to microscopic interaction mechanisms. These, along with other approaches (such as active particle theory or redistribution operators), allow one to derive, and thus mathematically justify, several models already known in the literature, while providing broader insight into their behavior. In this direction, future work will aim at deriving the reactive terms of [9] within a predator–prey framework.
Finally, as argued above, further attention will be devoted to analyze the behavior of the densities and the possible presence of interacting modes.

### v2

@@ 1 Introduction
Reaction-diffusion equations play a crucial role in various areas of applied sciences, offering a powerful framework for modeling diverse spatio-temporal dynamics in complex systems, that involve both interactions among the components and diffusive phenomena. Since the pioneering work of Alan Turing in 1952 [38], who first demonstrated how diffusion could destabilize a uniform equilibrium and lead to the emergence of stationary patterns, many reaction–diffusion models have been formulated at the macroscopic level by specifying partial differential equations for the concentrations or densities of interacting species.
Diffusion coefficients and reaction rates are introduced or fitted through phenomenological considerations. A well-known example is the Brusselator model, a classical autocatalytic reaction–diffusion system developed by Prigogine and collaborators [22, 33]. It has been extensively studied for its capacity to generate stable spatial patterns and to exhibit Turing instability under suitable parameter regimes [16, 32]. Moving forward, these models have been applied to a wide range of phenomena across disciplines. Examples include morphogenesis in autocatalytic chemical reactions [16, 32], vegetation patterns in arid landscapes [1, 14], epidemic spread [2, 15], lesion formation in autoimmune dynamics [6, 17], and even socioeconomic processes such as market dynamics and migration [3].
Beyond their broad applicability, classical reaction–diffusion systems possess a well developed mathematical theory explaining how patterns emerge from uniform states. In the two-species activator–inhibitor framework introduced by Turing [38] and extensively applied to biological systems in [30], diffusion-driven instability occurs when a homogeneous steady state, stable in the absence of diffusion, becomes unstable once diffusion is included, typically requiring a strong disparity between diffusion coefficients. Near the instability threshold, the dominant spatial modes select characteristic wavelengths, and in two-dimensional domains these modes can interact to produce a variety of stationary structures, including stripes, spots, labyrinthine patterns, and hexagonal arrays, as demonstrated numerically in early works like [29] and experimentally in chemical systems (e.g., [40]).
While one-dimensional settings allow only simple periodic configurations, two dimensions enable mode resonances, symmetry-breaking interactions, and the coexistence or competition of multiple pattern analyzed in depth through bifurcation and weakly nonlinear theory [20]. Weakly nonlinear (amplitude-equation) analysis has been a central tool for understanding this pattern-selection mechanism, for predicting the stability of different morphologies, and for developing more systematic criteria for spot–stripe selection and related transitions, as explored for instance in [25].
These classical approaches have proven highly successful in reproducing observed patterns across many domains, nevertheless, they often do not account for the underlying microscopic interactions between individual particles, agents, or molecules. As a result, the macroscopic parameters, such as diffusion coefficients and reaction constants, may lack a clear mechanistic interpretation. This limits the model’s predictive power, particularly in contexts where one seeks to relate observed macroscopic phenomena to their underlying physical or chemical processes.
In contrast, one can formulate macroscopic reaction–diffusion models directly from a microscopic description. This strategy allows the fundamental microscopic phenomena of interest to be modeled explicitly, with the macroscopic equations arising only afterward as collective effects of the microscopic dynamics. In doing so, the need to assign empirical parameter values is reduced or eliminated; instead, the resulting macroscopic expressions are grounded in the underlying microscopic physics of the problem.
To this aim, extended kinetic theory offers a powerful multiscale methodology, which has been employed in a variety of physical, chemical, and biological problems through the years (see for example [4, 12, 28, 37]). By starting from a mesoscopic description of particle interactions, it is possible to rigorously derive macroscopic reaction–diffusion systems through suitable hydrodynamic limits. These tools have been widely used in different frameworks, from classical Boltzmann theory of gas dynamics [5, 8, 21] to the models of cells and tissues (see e.g. [11] and references therein). Similarly, the micro-macro connection has been investigated in the case when the reactive collision operators are phenomenological models, such as Fokker-Planck or BGK [34, 35, 36], or for discrete velocity models [41].
@@
This approach not only provides explicit expressions for macroscopic parameters in terms of microscopic quantities but also reveals how collective behavior emerges from individual interactions. For the sake of completeness, it is important to remark that, from a broader perspective, kinetic theory offers the possibility of working across multiple scales, ultimately leading to a deeper understanding of collective phenomena and allowing, for example, insight into how a population is distributed around the mean value. Some steps in this direction have recently been taken in [10, 26].
In this paper, we focus on two such reaction-diffusion systems [8, 27], recently derived from kinetic models of gas mixtures involving monatomic and polyatomic species in proper multiple scale regimes. These models consider a mixture of interacting gases, initially modeled at the microscopic level using kinetic equations. From the microscopic formulation, the authors subsequently obtain a macroscopic formulation in the form of a reaction-diffusion system of equations. Unlike classical reaction–diffusion systems where diffusion and reaction rates are free parameters, the models derived in [8, 27] inherit nonlinear and cross-diffusion structures that arise directly from collision dynamics at the microscopic scale, leading to constraints and couplings absent in phenomenological models.
@@
Specifically, these references show how linear and nonlinear diffusion terms can be obtained through a rigorous limit of Boltzmann-like kinetic equations. More precisely, in the continuum limit, in [8] components densities are governed by a Brusselator-type description, where a further parameter appears in addition to those present in the classic model for autocatalytic reactions; whereas, in [27] a model with nonlinear cross diffusion is derived, reproducing the diffusive terms of [13] but differing from it in the reactive contributions. In both cases, the parameters in the macroscopic setting can be expressed in terms of the microscopic quantities, such as mass, internal energy and collision frequency , allowing for a physically grounded parameter selection and a deeper understanding of pattern formation mechanisms.
@@
The paper is organized as follows: we briefly recall the derivation of the two models as diffusive limit of proper kinetic equations in Section 2. The Turing instability and the pattern formation for the Brusselator-type model and the description with cross diffusion are discussed in Section 3 and 4, respectively: the emergence of non-uniform stable solutions is investigated through weakly nonlinear analysis, and confirmed and extended numerically. Some concluding remarks are given in Section 5.
@@ 2 Kinetic equations and diffusive limits
We first provide some preliminaries on the physical modeling adopted here to describe interacting mixtures of polyatomic species [18, 19]. Each gas species, say $C^{i}$, may have a number $l^{i}$ of possible discrete values for the internal energy. We denote each one of these values by $E^{i}_{j}$, $j=1,\ldots,l^{i}$. As a consequence, we treat the $i$-th species as composed by $l^{i}$ components $C^{i}_{j}$, $j=1,\ldots,l^{i}$, being the component $C^{i}_{j}$ characterized by the internal energy $E^{i}_{j}$. We report here a general encounter expressed as follows
$C^{i}_{j}+C^{h}_{k}\leftrightarrows C^{l}_{m}+C^{n}_{p}.$ (2.1)
In the case of mechanical collisions, in which molecules do not change their nature, we may have elastic scattering, when there is not energy dissipation due to absorption by the particles, and in the formulation (2.1) we have $(i,j)=(l,m)$ and $(h,k)=(n,p)$. If, instead, there is a variation of the internal energies of the pair of impinging particles, we have the collision (2.1) with $i=l$ and $h=n$, but $j\not=m$ and/or $k\not=p$. Finally, in case of chemical encounters, there is also a mass transfer and particles change the gas species they belong to. When this happens we have collision (2.1) with $(i,h)\neq(l,n)$ and $i\neq n$.
Indicating by $f_{j}^{i}(t,\mathbf{x},\mathbf{v})$ the distribution function associated to the component $C_{j}^{i}$ depending on time $t$, position $\mathbf{x}$, and molecular velocity $\mathbf{v}$, the Boltzmann-type operator for encounter (2.1) takes the following form (dependence on time and space is omitted)
$\mathcal{Q}_{i}^{XX}(f_{j}^{i},f_{k}^{h},f_{m}^{l},f_{p}^{n})=\int_{\mathbb{R}^{3}}\int_{\mathbb{S}^{2}}$ $\mathcal{H}(g^{2}-\delta_{ih}^{ln})g\sigma(g,\hat{\boldsymbol{\Omega}}\cdot\hat{\boldsymbol{\Omega}}^{\prime})$ (2.2)
$\times\left[\left(\dfrac{m_{i}m_{h}}{m_{l}m_{n}}\right)^{3}f_{m}^{l}(\mathbf{v}^{\prime})f_{p}^{n}(\mathbf{w}^{\prime})-f_{j}^{i}(\mathbf{v})f_{k}^{h}(\mathbf{w})\right]d\mathbf{w}d\hat{\boldsymbol{\Omega}}^{\prime}\,$,
where
    $(\mathbf{v},\mathbf{w})$ are the ingoing microscopic velocities of particles with mass $m_{i}$ and $m_{h}$, respectively;
    $(\mathbf{v}^{\prime},\mathbf{w}^{\prime})$ are the post interaction microscopic velocities of particles having mass $m_{l}$ and $m_{n}$ ($l=i$ and $n=h$ in mechanical encounters);
    the heaviside function $\mathcal{H}$ depends on the modulus of relative velocity $\mathbf{g}=\mathbf{v}-\mathbf{w}$ and on the energy gap $\delta_{ih}^{ln}=2m_{i}\,m_{h}\,(E_{m}^{l}+E_{p}^{n}-E_{j}^{i}-E_{k}^{h})/(m_{i}+m_{h})$ (the gap is null in elastic collisions);
    the cross sections $\sigma$ depends on the modulus of the ingoing relative velocity and on the deflection angle $\hat{\boldsymbol{\Omega}}\cdot\hat{\boldsymbol{\Omega}}^{\prime}$, being $\hat{\boldsymbol{\Omega}}$ the direction of vector $\mathbf{g}$.
The gaseous mixture considered in in [8, 27] is composed by two constituents diffusing in a host medium. The first constituent $Y$ is assumed to be polyatomic with two different energy levels $E_{1}$ and $E_{2}$; therefore, it can be split in two components $Y_{1}$ and $Y_{2}$. The second gas $Z$ is supposed to be monatomic, endowed with a unique energy level $E_{Z}$. We indicate by $m_{Y}$ and $m_{Z}$ the masses of polyatomic and monatomic gas particles, respectively.
@@ 2.1 Linear diffusion
It is important to underline that, through the derivation, an explicit expression of $n_{Z}$ as quadratic function of $n_{1}$ is deduced; this algebraic relation allows to reduce the number of variables in the diffusive limit.
In the following, the stability properties will be investigated in a regular open domain; if we suppose that the distribution functions $f_{i}^{\varepsilon}$ satisfy given initial conditions at time $t=0$ and specular reflection conditions at the boundary, then, as proved in [5], the macroscopic densities exactly obeys equations (2.13), with proper initial data and no-flux conditions at the boundary.
The procedure briefly outlined here shows that the coefficients of the macroscopic observables are positive constants that can be explicitly expressed in terms of the underlying microscopic parameters, such as collision frequencies, particle masses, and discrete energy levels. In this way, the collective behavior encoded in the macroscopic equations remains directly related to the interaction mechanisms governing the particles dynamics at the microscopic scale. This correspondence not only enhances the physical meaning of the macroscopic coefficients but also introduces natural constraints on their admissible values, thereby delimiting the range of validity of the resulting continuum model.
@@ 2.2 Nonlinear cross diffusion
where $N=n_{1}+n_{2}$ is the total number density of the polyatomic component and the use of an algebraic relation relating $n_{1}$, $n_{2}$ and $n_{Z}$ has been made ; as shown in the previous subsection, equations are coupled with proper initial and boundary conditions. Once again, the parameters are positive constants, depending on microscopic parameters as provided below
@@ 3.1 Turing instability
By expressing the coefficients in terms of the microscopic parameters of the mixture through (2.14)-(2.17) and, for illustrative purposes, without reference to any specific physical scenarios, assigning the following values
@@
we identify the values for energy levels $E_{2}$ and $E_{Z}$ leading to condition (3.15) and (3.16) to be satisfied, as reported in Figure 1. In particular, we can distinguish four regions. In region I, condition (3.3) is not satisfied and the spatially homogeneous equilibrium is unstable. In regions II, III, and IV, instead, (3.3) holds, and either both conditions (3.15) and (3.16), only (3.16), or neither of the conditions in (3.15) and (3.16) are satisfied, respectively. Consequently, pattern formation can arise only when taking parameters in region II. It should be emphasized that, within region II, the parameter $d$ is not restricted to the value 1, which once again highlights the broader spectrum of scenarios due to the presence of this additional parameter.
@@ 3.2 Weakly nonlinear analysis
where by c.c. we indicate the complex conjugate of the first term in (3.34) and $\sigma=\dfrac{a\xi\,\sqrt{d}\,}{1+a\xi\,\sqrt{d}\,}$.
@@
$+\sum_{\begin{subarray}{c}j=1,2,3\\ l\equiv j+1\mbox{ \tiny mod }3\end{subarray}}\left(\begin{array}[]{c}X_{1}\\ Y_{1}\end{array}\right)\,\mathcal{W}_{j}\,\overline{\mathcal{W}}_{l}\,e^{\,i\,(\mathbf{k}_{j}-\mathbf{k}_{l})\cdot\mathbf{x}}+\mbox{ c.c.}\,$,
where here c.c. collect the complex conjugates of the explicit terms on the right-hand side.
@@ 3.3 Numerical simulations
In this section we investigate the behavior of system (2.13) by performing numerical simulations. We simulate a two-dimensional square domain of length $L=12\pi$, discretized into a $\times100$ grid. The solution is initialized by adding a random perturbation to the equilibrium values. The diffusive terms are discretized using a second-order finite-difference method, the source terms are evaluated locally, and the solution is marched in time from the initial conditions until convergence using a first-order forward-Euler scheme. It is worth noting that once the solution is converged to a steady state, the error associated with the time discretization vanishes. Therefore, using a first-order time-integration scheme is acceptable for our purposes. The time step is dynamically adpated throughout the simulation to maintain numerical stability by keeping the local von Neumann number - determined from the maximum diffusion coefficient in the domain - under a given threshold. At the boundary of the domain, we impose zero-gradient Neumann conditions. We do so by utilizing a layer of ghost cells and adapting their value as the simulation evolves. The method is implemented in a modified version of the open-source Hyper2D solver [9] tailored to reaction-diffusion systems.
@@
For the sake of brevity, we do not report the patterns in regions I and II, since their shape is similar to that of the patterns in regions IV and V, respectively; the only difference concerns the phase angle $\phi$, which passes from $\pi$ to 0, giving rise to the so-called reentrant hexagons [39]. Although our model includes an additional parameter $d$, it reproduces qualitatively the same types of patterns reported in the literature [32], albeit with some quantitative differences. Different parameter values, as well as different choices of the initial datum, lead to different patterns; the different ways in which instabilities interact with nonlinearities can sometimes result in the coexistence of multiple pattern types, as illustrated in Figure 4.
@@ 4.1 Turing instability and weakly nonlinear analysis
As in the previous case, we want to focus on the microscopic features of the model, by finding suitable sets of the microscopic parameters that lead to the conditions above to be satisfied by the macroscopic coefficients, through the relations (2.21)-(2.23). For our purposes, we assign the following values as an illustrative example, without referring to any specific physical scenarios
@@ 4.2 Numerical simulations
Although some patterns presented here may resemble those reported in other contexts, it is worth underlying that the results shown are entirely new; as already noted, the proposed model reproduces the diffusive mechanism of [13], but its reactive terms are substantially different.
Moreover, we point out that, as shown also in the one-dimensional case in [27], the numerical results suggest that the profile of $N$ is sum of more than one mode in each direction $\mathbf{k}_{i}$. Therefore, a deeper analysis of the solution near the critical threshold is needed; this will be subject of future investigation.
@@ 5 Conclusions
The proposed analysis shows the strength of these tools in linking macroscopic dynamics to microscopic interaction mechanisms. These, along with other approaches (such as active particle theory or redistribution operators), allow one to derive, and thus mathematically justify, several models already known in the literature, while providing broader insight into their behavior. In this perspective, future work will be devoted to the derivation of alternative reactive terms, such as those proposed in [13]. A careful calibration of these terms, based also on a deep knowledge of the underlying microscopic interaction mechanisms, is essential, as the intrinsic nonlinearities in them play a crucial role in enhancing the realism of the models. These nonlinear effects not only allow a more faithful representation of the processes but also provide the key element to reproducing and interpreting the rich variety of spatial patterns observed in real systems.
Finally, as argued above, further attention will be devoted to analyze the behavior of the densities and the possible presence of interacting modes whose study is particularly useful in applications because the coupling among spatial modes can lead to the emergence of complex patterns, influene their stability, and affect their shape. Understanding these interactions could help to better clarify the mechanisms behind complex structures observed in real ecosystems, such as mixed and oscillating patterns, and to interpret empirical data and evidence. In addition, in this work, we have focused on the patterns predicted analytically near the critical thresholds, showing that stable spatial structures emerge from perturbations of the equilibrium. More complex and temporally oscillating patterns may arise from initial conditions far from these thresholds [7], and we plan to investigate these richer dynamics in future work.